# Mean-field theory of 1+1D $\mathbb{Z}_{2}$ lattice gauge theory with matter

### v1

@@ I Introduction
Here we study a 1+1D $\mathbb{Z}_{2}$ LGT where dynamical charges are coupled to gauge fields, where we generalise the problem by including superconducting (SC) terms which explicitly break the U$(1)$ conservation of the charge number. We consider the so called physical sector without background charges which is defined by the set of local generators of the gauge structure, which is the LGT counterpart of the Gauss law [4]. Individual particles (partons) bind into dimers (mesons) in the presence of the $\mathbb{Z}_{2}$ electric field term in the Hamiltonian, which induces dynamics in the $\mathbb{Z}_{2}$ gauge field [24]. The confinement mechanism can be understood by mapping the system to the so called non-local string length basis, where confinement can be related to the breaking of the translational symmetry [25]. In addition, a smooth confinement-deconfinement crossover was uncovered as a function of temperature [26].
@@
Figure 1: One-dimensional $\mathbb{Z}_{2}$ lattice gauge theory with superconducting term Eq. (1). (a) The Gauss law constraint Eq. (2) ensures that the $\mathbb{Z}_{2}$ electric field changes its prefactor across an occupied lattice site, which allows us to define the $\mathbb{Z}_{2}$ strings (orange lines) labeling $\tau^{x}=-1$ and anti-stings (no line) which labels $\tau^{x}=+1$. (b) Different regimes of the 1+1D $\mathbb{Z}_{2}$ LGT Eq. (1). In the first row (from above) we sketch the free parton regime, in the second row we sketch the regime where partons are confined into mesons, in the third row we illustrate deconfined partons with SC term, where the U(1) symmetry is explicitly broken (meson creation and annihilation), and in the last row we illustrate the confined regime with the SC term. (c) A qualitative sketch of a phase diagram in the deconfined regime $h=0$, which exhibits a deconfined parton Luttinger liquid (LL) on the $\lambda=0$ line, a deconfined symmetry protected state (SPT) for intermediate fillings which correspond to $t<2|\mu|$, ferromagnetic (FM) symmetry broken phase at low filling, and antiferromagnetic (AFM) symmetry broken phase at high filling. (d) A qualitative sketch of a phase diagram in the regime $h\neq 0$, which exhibits a confined meson LL on the $\lambda=0$ line, confined Higgs phase for $\lambda\neq 0$ up to moderate doping, and a symmetry broken AFM phase for high doping.
@@
In our work we first establish the phase diagram of the 1+1D $\mathbb{Z}_{2}$ LGT as a function of filling and the strength $\lambda$ of the SC pairing term. We considering the regime with and without the confining $\mathbb{Z}_{2}$ electric field term $h$ and reproduce the phase diagram of Ref. [28]. However, our analysis focuses on the microscopic picture of the doped $\mathbb{Z}_{2}$ LGTs; in particular we study the confinement in the presence of the SC term. In addition to entanglement entropy calculations, we also study the $\mathbb{Z}_{2}$ invariant Green’s function and sample string-length histograms in order to probe the confinement. We show that confining features are also present in the topological trivial states at finite filling, where the $\mathbb{Z}_{2}$ electric field term strength is zero, $h=0$. Notably, this establishes the symmetry-protected topological (SPT) transition in the Kitaev chain as a confinement transition, akin to the topological confinement transition known from the $2+1$D perturbed toric code [29, 30, 31, 32].
@@
We then derive a mean-field description of the $\mathbb{Z}_{2}$ LGT by effectively decoupling matter and $\mathbb{Z}_{2}$ fields, but enforcing the Gauss law on the mean-field level. We find that the mean-field analysis qualitatively captures the main features of the $\mathbb{Z}_{2}$ LGT aside from the U$(1)$ symmetric critical line, since the mean-field theory explicitly breaks this U$(1)$ symmetry. In addition, we probe the confinement in the mean-field theory and reveal qualitative agreement with the original $\mathbb{Z}_{2}$ LGT. Finally, we explicitly compare the ground state energies and the $\mathbb{Z}_{2}$ electric field polarization in the original $\mathbb{Z}_{2}$ LGT and its mean-field theory, where we find excellent agreement between both theories.
@@ II Model
^$\hat{\mathcal{H}}=-t\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}_{j}+\text{h.c.}\right)-h\sum_{{\langle i,j\rangle}}\hat{\tau}^{x}_{{\langle i,j\rangle}}\\ +\lambda\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\hat{n}_{j}.$ (1)
@@
which define the Gauss law of the $\mathbb{Z}_{2}$ lattice gauge theory. These operators commute on different lattice sites $[\hat{G}_{i},\hat{G}_{j}]=0$ and with the Hamiltonian $[\hat{G}_{j},\hat{\mathcal{H}}]=0$ [4]. The eigenvalues of Eq. (2) can take two possible values $g_{j}=\pm 1$ and we constrain our Hilbert space to the sub-sector where the eigenvalues are equal to $g_{j}=+1,\forall j$. The $\mathbb{Z}_{2}$ electric field is thus anti-aligned across an occupied lattice site and aligned across an empty site, see Fig. 1(a) and (b). Such constraint allows us to define $\mathbb{Z}_{2}$ electric strings and anti-strings which connect the pairs of individual particles and reflect the orientation of the electric field $\tau^{x}_{{\langle i,j\rangle}}$ [24, 25]. We define the string as $\tau^{x}=-1$ and the anti-string as $\tau^{x}=+1$, see Fig.1(a).
@@
[…]
and can be regarded as a $\mathbb{Z}_{2}$ invariant superconducting (SC) term. This term explicitly breaks the U$(1)$ symmetry of the charges and thus the particle number conservation, since it adds or removes a pair of particles while conserving the Gauss law constraint.
The $\mathbb{Z}_{2}$ electric field term with strength $h$ induces dynamics in the gauge fields and leads to a linear confining potential for the strings [24, 25]. It has been show that the $\mathbb{Z}_{2}$ lattice gauge theory in Eq. (1) without superconducting terms ($\lambda=0$) exhibits confinement of particles into mesons for any nonzero value of the $\mathbb{Z}_{2}$ electric field term $h$ [24, 25]. Confined mesons remain mobile and form a meson Luttinger liquid (LL), which differs from the $h=\lambda=0$ case where a LL of the individual partons ($\hat{a}$) forms [24, 25, 27]. The distinct nature of the two LLs at $h=0$ and $h\neq 0$ (both at $\lambda=0$) can be observed in the doubling of the period of Friedel oscillations in the confined regime [24]. A general solution of the confinement problem in the 1+1D $\mathbb{Z}_{2}$ LGT without superconducting term was found by mapping the above model to a non-local string-length basis, where it has been shown that the translational symmetry breaking in the new basis is directly related to the confinement in the original $\mathbb{Z}_{2}$ LGT [25]. Furthermore, it has been shown that the model features a smooth confinement-deconfinement crossover as a function of temperature [26].
@@ III.2.1 Value of the entanglement entropy
To determine the phase diagram of the model, we first consider the entanglement entropy $S(x)$ as a function of filling $n$, and the strength of the SC term $\lambda$ at different $\mathbb{Z}_{2}$ electric field strengths $h$. To calculate the entanglement entropy we cut our chain into two subsystems $A$ and $B$, where we denote the site of the cut with $x$. The length of the subsystem $A$ is thus $x$ and the length of the subsystem $B$ is $L+1-x$. Note that the total spin system contains $L+1$ spins, since we simulate the gauge fields directly. The cuts are thus between links $\tau_{x-1,x}$ and $\tau_{x,x+1}$ , i.e, on the matter site $x$.
Figure 2: Entanglement entropy of the $\mathbb{Z}_{2}$ LGT Eq. (1), after integrating out matter fields, as a function of filling $n$ and value of the SC term $\lambda$. (a) Symmetric behavior $(n=0.5+\Delta n)\leftrightarrow(n=0.5-\Delta n)$ is observed for $h=0$. The entanglement entropy is larger for positive values of the SC term $\lambda>0$ than for their negative values. We notice a trapezoid shape where the entanglement entropy drastically decreases at low filling $n\lesssim 0.15$ and high filling $n\gtrsim 0.85$, which corresponds to the transition between the topological phase and trivial phase in the Kitaev chain. (b) The entanglement entropy gradually increases with increasing filling $n$ at finite field term $h/t=1.0$ The white dots represent the data points, obtained from DMRG, from which the software triangulates the heat map values.
@@
Another interesting feature in Fig. 2 is the clear distinction between $\lambda>0$ and $\lambda<0$, which can be seen in both diagrams. We observe a large plateau of substantial entanglement entropy in the absence of the $\mathbb{Z}_{2}$ electric field term ($h=0$) for $\lambda>0$, see Fig. 2(a). This plateau narrows for stronger values of $\lambda$ indicating that there is a phase transition as a function of filling at very high and low fillings. In contrast, there is a large plateau of almost zero entanglement entropy for $\lambda<0$. We observe some non-zero entanglement entropy features again at low and high fillings. These features again narrow in a similar way as the plateau for $\lambda>0$ which indicates that the same phase transition occurs also for $\lambda<0$. Indeed, by a gauge transformation $\hat{a}_{j}\rightarrow e^{i\frac{\pi}{2}}\hat{a}_{j}$, the model Eq. (1) is equivalent at ±$\pm\lambda$, but this gauge transformation has a non-trivial effect on the entanglement we calculate after integrating out matter fields, i.e., directly in the $\hat{\tau}^{x}_{{\langle i,j\rangle}}$ basis, see Appendix B.
@@ III.2.2 Central charge
$S(x)=S_{0}+\frac{c}{6}\log{\left[\left(\frac{2L^{\prime}}{\pi}\right)\sin{\left(\frac{\pi x}{L^{\prime}}\right)}\right]}.$ ) . (4)
@@
As already stated earlier we can eliminate the gauge fields in the regime where $h=0$ by attaching a string of gauge fields to our charge operators [4, 24]. This mapping gives us the celebrated Kitaev model for a 1D superconductor [35], see also Appendix A. Such model is known to exhibit a topological non-trivial phase for $|\mu|<2|t|$ [35, 37], which corresponds to a finite range of fillings $10<n_{c1}<n<n_{c2}<$ 1. By plotting the same results as the function of chemical potential $\mu$, we confirm that the lines in Fig. 3(a) indeed correspond to the $|\mu|=2t$ border between the topological trivial and non-trivial phase, see Appendix E.
@@ III.3.1 Green’s function
In the absence of the SC terms it decays as a power-law in the deconfined regime $h=0$ and exponentially in the confined regime $h\neq 0$. We note that $\mathcal{G}(x)$ can be viewed as a one-dimensional version of the Fredenhagen-Marcu order parameter considered in higher dimensions [48]. In a deconfined phase with free partons (confined phase with partons bound into mesons) the Green’s function decays with a power-law (exponentially) in $1+1$D dimensions.
@@ III.3.2 String length histograms
We note that for non-zero SC term, $\lambda/t=-1$, the string and anti-string length distributions differ for low fillings $n\lesssim 0.2$, where the string-length peak at $\ell=1$ is much higher while the anti-string length distribution is significantly broader, see Fig. 5(c). This confirms the conclusion from our Green’s function analysis in Fig. 4(c) that there exists a spontaneously confined phase for low $n\lesssim 0.2$ and high $n\gtrsim 0.8$ fillings when the SC term is non-zero $\lambda\neq 0$ and $h=0$. The microscopic picture can also be easily described: At such low fillings all particles come from pair fluctuations generated by the SC terms. All particles thus come in pairs, i.e. in forms of mesons. The mesons are then soon annihilated, on average much sooner than the parton would be able to hop, hence they become confined.
In the spin model, which we simulate with DMRG, the picture is even more clear. For simplicity we can consider the case when $\lambda=-t$, where the model becomes a simple transverse field Ising model, see Appendix B. There, mesons are simple excitations in a form of flipped spins in the ferromagnetic state realized at $\mu<2$ (i.e. $n\leq 0.2$) and $h=0$, see also the discussion in Sec. IV.
@@
Figure 6: Sketches of phase diagrams of the $\mathbb{Z}_{2}$ LGT as a function of SC term $\lambda$ and $\mathbb{Z}_{2}$ electric field term $h$, for increasing matter densities. (a) Free partons form a Luttinger liquid when $h=0$ and $\lambda=0$, which turns into meson LL for any non-zero value of $h\neq 0$ and if $\lambda=0$. On the other hand the $\lambda\neq 0$ term explicitly breaks the U(1) symmetry and the parton LL becomes a deconfined SPT at $h=0$, which at low filling has a transition point to a confined FM state, with spontaneous symmetry breaking (SSB). For any non-zero $h\neq 0$ and $\lambda\neq 0$, the system is in a confined Higgs state. (b) At intermediate fillings the phase diagram is similar to (a) with a difference that the system remains in a deconfined SPT phase at $h=0$, and doesn’t transition to a confined state. (c) At high filling a similar transition as in (a) occurs as a function of $\lambda$ for $h=0$. The only difference is that in this regime the SSB state is an AFM state and that it remains stable also for $h\neq 0$. (d) For comparison we also present the phase-diagram of the $2+1$D $\mathbb{Z}_{2}$ LGT as proposed by Fradkin and Shenker [30].

### v2

@@ I Introduction
Here we study a 1+1D $\mathbb{Z}_{2}$ LGT where dynamical charges are coupled to gauge fields, where we generalise the problem by including superconducting (SC) terms which explicitly break the U(1)(1) conservation of the charge number. We consider the so called physical sector without background charges which is defined by the set of local generators of the gauge structure, which is the LGT counterpart of the Gauss law [4]. Individual particles (partons) bind into dimers (mesons) in the presence of the $\mathbb{Z}_{2}$ electric field term in the Hamiltonian, which induces dynamics in the $\mathbb{Z}_{2}$ gauge field [24]. The confinement mechanism can be understood by mapping the system to the so called non-local string length basis, where confinement can be related to the breaking of the translational symmetry [25]. In addition, a smooth confinement-deconfinement crossover was uncovered as a function of temperature [26].
@@
Figure 1: One-dimensional $\mathbb{Z}_{2}$ lattice gauge theory with superconducting term Eq. (1). (a) The Gauss law constraint Eq. (2) ensures that the $\mathbb{Z}_{2}$ electric field changes its prefactor across an occupied lattice site, which allows us to define the $\mathbb{Z}_{2}$ strings (orange lines) labeling $\tau^{x}=-1$ and anti-strings (no line) which labels $\tau^{x}=+1$. (b) Different regimes of the 1+1D $\mathbb{Z}_{2}$ LGT Eq. (1). In the first row (from above) we sketch the free parton regime, in the second row we sketch the regime where partons are confined into mesons, in the third row we illustrate deconfined partons with SC term, where the U(1) symmetry is explicitly broken (meson creation and annihilation), and in the last row we illustrate the confined regime with the SC term. (c) A qualitative sketch of a phase diagram in the deconfined regime $h=0$, which exhibits a deconfined parton Luttinger liquid ($L$) on the $\lambda=0$ line, a deconfined symmetry protected state (SPT) for intermediate fillings which correspond to $t<2|\mu|$, ferromagnetic (FM) symmetry broken phase at low filling, and antiferromagnetic (AFM) symmetry broken phase at high filling. (d) A qualitative sketch of a phase diagram in the regime $h\neq 0$, which exhibits a confined meson LL on the $\lambda=0$ line, confined Higgs phase for $\lambda\neq 0$ up to moderate doping, and a symmetry broken AFM phase for high doping.
@@
In our work we first establish the phase diagram of the 1+1D $\mathbb{Z}_{2}$ LGT as a function of filling and the strength $\lambda$ of the SC pairing term. We considering the regime with and without the confining $\mathbb{Z}_{2}$ electric field term $h$ and reproduce the phase diagram of Ref. [28]. However, our analysis focuses on the microscopic picture of the doped $\mathbb{Z}_{2}$ LGTs; in particular we study the confinement in the presence of the SC term. In addition to entanglement entropy calculations, we also study the $\mathbb{Z}_{2}$ invariant Green’s function and sample string-length histograms in order to probe the confinement. We show that confining features are also present in the topological trivial states at finite filling, where the $\mathbb{Z}_{2}$ electric field term strength is zero, $h=0$. Notably, this establishes the symmetry-protected topological (SPT) transition in the Kitaev chain as a confinement transition, akin to the topological confinement transition known from the 2+12+1D perturbed toric code [29, 30, 31, 32].
@@
We then derive a mean-field description of the $\mathbb{Z}_{2}$ LGT by effectively decoupling matter and $\mathbb{Z}_{2}$ fields, but enforcing the Gauss law on the mean-field level. We find that the mean-field analysis qualitatively captures the main features of the $\mathbb{Z}_{2}$ LGT aside from the U(1)(1) symmetric critical line, since the mean-field theory explicitly breaks this U(1)(1) symmetry. In addition, we probe the confinement in the mean-field theory and reveal qualitative agreement with the original $\mathbb{Z}_{2}$ LGT. Finally, we explicitly compare the ground state energies and the $\mathbb{Z}_{2}$ electric field polarization in the original $\mathbb{Z}_{2}$ LGT and its mean-field theory, where we find excellent agreement between both theories.
@@ II Model
$\hat{\mathcal{H}}=-t\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}_{j}+\text{h.c.}\right)-h\sum_{{\langle i,j\rangle}}\hat{\tau}^{x}_{{\langle i,j\rangle}}\\ +\lambda\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\hat{n}_{j}.$ (1)
@@
which define the Gauss law of the $\mathbb{Z}_{2}$ lattice gauge theory. These operators commute on different lattice sites $\hat{G}_{i},\hat{G}_{j}]=0$ and with the Hamiltonian $\hat{G}_{j},\hat{\mathcal{H}}]=0 [$4]. The eigenvalues of Eq. (2) can take two possible values $g_{j}=\pm 1$ and we constrain our Hilbert space to the sub-sector where the eigenvalues are equal to $g_{j}=+1,\forall j$. The $\mathbb{Z}_{2}$ electric field is thus anti-aligned across an occupied lattice site and aligned across an empty site, see Fig. 1(a) and (b). Such constraint allows us to define $\mathbb{Z}_{2}$ electric strings and anti-strings which connect the pairs of individual particles and reflect the orientation of the electric field $\tau^{x}_{{\langle i,j\rangle}}$ [24, 25]. We define the string as $\tau^{x}=-1$ and the anti-string as $\tau^{x}=+1$, see Fig.1(a).
@@
$\hat{\mathcal{H}}_{\lambda}=\lambda\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}^{\dagger}_{j}+\text{h.c.}\right),$ (3)
and can be regarded as a $\mathbb{Z}_{2}$ invariant superconducting (SC) term. This term explicitly breaks the U(1)(1) symmetry of the charges and thus the particle number conservation, since it adds or removes a pair of particles while conserving the Gauss law constraint.
The $\mathbb{Z}_{2}$ electric field term with strength $h$ induces dynamics in the gauge fields and leads to a linear confining potential for the strings [24, 25]. This term explicitly breaks the global $\mathbb{Z}_{2}$ magnetic symmetry for $h\neq 0$ [28]. It has been show that the $\mathbb{Z}_{2}$ lattice gauge theory in Eq. (1) without superconducting terms ($\lambda=0$) exhibits confinement of particles into mesons for any nonzero value of the $\mathbb{Z}_{2}$ electric field term $h$ [24, 25]. Confined mesons remain mobile and form a meson Luttinger liquid (LL), which differs from the $h=\lambda=0$ case where a LL of the individual partons ($\hat{a}$) forms [24, 25, 27]. The distinct nature of the two LLs at $h=0$ and $h\neq 0$ (both at $\lambda=0$) can be observed in the doubling of the period of Friedel oscillations in the confined regime [24]. A general solution of the confinement problem in the 1+1D $\mathbb{Z}_{2}$ LGT without superconducting term was found by mapping the above model to a non-local string-length basis, where it has been shown that the translational symmetry breaking in the new basis is directly related to the confinement in the original $\mathbb{Z}_{2}$ LGT [25]. Furthermore, it has been shown that the model features a smooth confinement-deconfinement crossover as a function of temperature [26].
@@ III.2.1 Value of the entanglement entropy
To determine the phase diagram of the model, we first consider the bipartite entanglement entropy $S(x)$ as a function of filling $n$, and the strength of the SC term $\lambda$ at different $\mathbb{Z}_{2}$ electric field term strengths $h$. To calculate the entanglement entropy we cut our chain into two subsystems $A$ and $B$, where we denote the site of the cut with $x$. The length of the subsystem $A$ is thus $x$ and the length of the subsystem $B$ is L+1xL+1-x. Note that the total spin system contains $L+1$ spins, since we simulate the gauge fields directly. The cuts are thus between links $\tau_{x-1,x}$ and $\tau_{x,x+1}$ , i.e, on the matter site $x$. The bipartite entanglement entropy $S(x)$ was extracted from the MPS directly from the singular value decomposition at a given bond $x$, which divides the system into subsystem $A$ and $B$ [33].
Figure 2: Entanglement entropy of the $\mathbb{Z}_{2}$ LGT Eq. (1), after integrating out matter fields, as a function of filling $n$ and value of the SC term $\lambda$. (a) Symmetric behavior $(n=0.5+\Delta n)\leftrightarrow(n=0.5-\Delta n)$ is observed for $h=0$. The entanglement entropy is larger for positive values of the SC term $\lambda>0$ than for their negative values. We notice a trapezoid shape where the entanglement entropy drastically decreases at low filling $n\lesssim 0.15$ and high filling $n\gtrsim 0.85$, which corresponds to the transition between the topological phase and trivial phase in the Kitaev chain. (b) The entanglement entropy gradually increases with increasing filling $n$ at finite field term $h/t=1.0$ . The white dots represent the data points, obtained from DMRG, from which the software triangulates the heat map values.
@@
Another interesting feature in Fig. 2 is the clear distinction between $\lambda>0$ and $\lambda<0$, which can be seen in both diagrams. We observe a large plateau of substantial entanglement entropy in the absence of the $\mathbb{Z}_{2}$ electric field term ($h=0$) for $\lambda>0$, see Fig. 2(a). This plateau narrows for stronger values of $\lambda$ indicating that there is a phase transition as a function of filling at very high and low fillings. In contrast, there is a large plateau of almost zero entanglement entropy for $\lambda<0$. We observe some non-zero entanglement entropy features again at low and high fillings. These features again narrow in a similar way as the plateau for $\lambda>0$ which indicates that the same phase transition occurs also for $\lambda<0$. Indeed, by a unitary transformation $\hat{a}_{j}\rightarrow e^{i\frac{\pi}{2}}\hat{a}_{j}$, the model Eq. (1) is equivalent at $\pm\lambda$, but this transformation has a non-trivial effect on the entanglement we calculate after integrating out matter fields, i.e., directly in the $\hat{\tau}^{x}_{{\langle i,j\rangle}}$ basis, see Appendix B.
@@ III.2.2 Central charge
[…]
@@
As already stated earlier we can eliminate the gauge fields in the regime where $h=0$ by attaching a string of gauge fields to our charge operators [4, 24]. This mapping gives us the celebrated Kitaev model for a 1D superconductor [35], see also Appendix A. Such model is known to exhibit a topological non-trivial phase for $|\mu|<2|t|$ [35, 37], which corresponds to a finite range of fillings $0<n_{c1}<n<n_{c2}<1$. By plotting the same results as the function of chemical potential $\mu$, we confirm that the lines in Fig. 3(a) indeed correspond to the $|\mu|=2t$ border between the topological trivial and non-trivial phase, see Appendix E.
@@ III.3.1 Green’s function
In the absence of the SC terms it decays as a power-law in the deconfined regime $h=0$ and exponentially in the confined regime $h\neq 0$. We note that $\mathcal{G}(x)$ can be viewed as a one-dimensional version of the Fredenhagen-Marcu order parameter considered in higher dimensions [48]. In a deconfined phase with free partons (confined phase with partons bound into mesons) the Green’s function decays with a power-law (exponentially) in 1+11+1D dimensions. We note that since we integrate out the matter by using the Gauss law and directly simulate the system by using the spin-1/2 model, we had to express the Green’s function in terms of the spin operators; see Eq. (16) and the Appendices B and D.
@@ III.3.2 String length histograms
We note that for non-zero SC term, $\lambda/t=-1$, the string and anti-string length distributions differ for low fillings $n\lesssim 0.2$, where the string-length peak at $\ell=1$ is much higher while the anti-string length distribution is significantly broader, see Fig. 5(c). This confirms the conclusion from our Green’s function analysis in Fig. 4(c) that there exists a spontaneously confined phase for low $n\lesssim 0.2$ and high $n\gtrsim 0.8$ fillings when the SC term is non-zero $\lambda\neq 0$ and $h=0$. The microscopic picture can also be easily described: At such low fillings all particles come from pair fluctuations generated by the SC terms. All particles thus come in pairs, i.e. in forms of mesons. The mesons are then soon annihilated, on average much sooner than the parton would be able to hop, hence they become confined. Similar pair-fluctuating mechanism is at work at high fillings, where the chain is almost completely filled, and the SC terms annihilate pairs, which are then again generated, before partons are able to hop.
In the spin model, which we simulate with DMRG, the picture is even more clear. For simplicity we can consider the case when $\lambda=-t$, where the model becomes a simple transverse field Ising model, see Appendix B. There, mesons are simple excitations in a form of flipped spins in the ferromagnetic state realized at $\mu<-2t$ (i.e. $n\leq 0.2$) and $h=0$, see also the discussion in Sec. IV.
@@
Figure 6: Sketches of phase diagrams of the $\mathbb{Z}_{2}$ LGT as a function of SC term $\lambda$ and $\mathbb{Z}_{2}$ electric field term $h$, for increasing matter densities. (a) Free partons form a Luttinger liquid when $h=0$ and $\lambda=0$, which turns into meson LL for any non-zero value of $h\neq 0$ and if $\lambda=0$. On the other hand the $\lambda\neq 0$ term explicitly breaks the U(1) symmetry and the parton LL becomes a deconfined SPT at $h=0$, which at low filling has a transition point to a confined FM state, with spontaneous symmetry breaking (SSB). For any non-zero $h\neq 0$ and $\lambda\neq 0$, the system is in a confined Higgs state. (b) At intermediate fillings the phase diagram is similar to (a) with a difference that the system remains in a deconfined SPT phase at $h=0$, and doesn’t transition to a confined state. (c) At high filling a similar transition as in (a) occurs as a function of $\lambda$ for $h=0$. The only difference is that in this regime the SSB state is an AFM state and that it remains stable also for $h\neq 0$. (d) For comparison we also present the phase-diagram of the 2+12+1D $\mathbb{Z}_{2}$ LGT as proposed by Fradkin and Shenker [30].